OpenStudy (anonymous):

Without using a calculator, find the matching value for f in each function. Give exact answers.

OpenStudy (anonymous):

\[f=\cos (-\frac{ 3\pi }{ 2 })\]

OpenStudy (anonymous):

find the point on the unit circle corresponding to the angle of \(-\frac{3\pi}{2}\) cosine will be the first coordinate

OpenStudy (anonymous):

thats 300 degrees and the first coordinate is 1/2 but thats not the correct answer...:/

OpenStudy (anonymous):



OpenStudy (anonymous):

oh oh oh i was looking at the coordinates of y

OpenStudy (anonymous):

so its 270 degrees.

OpenStudy (anonymous):

forget degrees. angles measure in radians you should be right at the top of the circle at \((0,1)\)

OpenStudy (anonymous):

At the top of my circle i have pi/2...

OpenStudy (anonymous):

yes that is true, but if you go around the circle in the negative direction (clockwise) then you end up at the same point

OpenStudy (anonymous):

but how does that change it? you end up at the same place...

OpenStudy (anonymous):

hope that is clear count in units of \(\frac{\pi}{2}\) going the other direction

OpenStudy (anonymous):

you do end up in the same place, so the cosine is 0 and the sine is 1

OpenStudy (anonymous):

in other words \(\cos(\frac{\pi}{2})=\cos(-\frac{3\pi}{2})\)

OpenStudy (anonymous):

both sine and cosine are periodic with period \(2\pi\)

OpenStudy (anonymous):

okay, okay i can see that cause if the bottom of the circle is 3pi/2 that the opposite of that would be negative. AKA the upper half of the circle. okay...

OpenStudy (anonymous):

or just count, either way

OpenStudy (anonymous):

okay it just kinda had to click in my brain.. so f=cos(-3pi/2) is the exact same thing as f=cos(pi/2)